OpenStudy (stuck-help):

can someone help me get he value of "X" just set up the equation @Natash18 @Lighthouseluver

OpenStudy (stuck-help):

@Natasha18

OpenStudy (whpalmer4):

midpoint of a line is just the average of the end points. If you have a line going from \((3,5)\) to \((7,11)\), the midpoint will be at \[(\frac{3+7}2,\frac{5+11}2) = (5,8)\]

OpenStudy (stuck-help):

so how do i do my question

OpenStudy (stuck-help):

@whpalmer4

OpenStudy (whpalmer4):

you need to find midpoint of line from point A to point C the x coordinate will be the average of the x-coordinate of A and the x coordinate of C the y coordinate will be the average of the y coordinate of A and the y coordinate of C to find the average of two numbers, add them, then divide the sum by 2

OpenStudy (natasha18):

A (0 ; 0) B (2 ; 2) C (5 ; - 1) Midpoint of [AC] xL = (xA + xC)/2 = (0 + 5)/2 = 5/2 yL = (yA + yC)/2 = (0 - 1)/2 = - 1/2 → L (5/2 ; - 1/2) Midpoint of [BC] xM = (xB + xC)/2 = (2 + 5)/2 = 7/2 yM = (yB + yC)/2 = (2 - 1)/2 = 1/2 → M (7/2 ; 1/2) Equation of the line (LM) The typical equation of a line is: y = mx + b → where m: slope and where b: y-intercept To calculate m m = (yM - yL) / (xM - xL) m = [(1/2) - (- 1/2)] / [(7/2) - (5/2)] m = [(1/2) + (1/2)] / [(2/2)] m = 1 / 1 m = 1 ← this is the slope of the line (LM) Equation of the line (AB) The typical equation of a line is: y = mx + b → where m: slope and where b: y-intercept To calculate m m = (yB - yA) / (xB - xA) m = (2 - 0) / (2 - 0) m = 1 ← this is the slope of the line (AB) → Two lines are parallel if they have the same slope. It's the case for (LM) and (AB). Distance LM xLM = xM - xL = (7/2) - (5/2) = 2/2 = 1 yLM = yM - yL = (1/2) - (- 1/2) = (1/2) + (1/2) = 1 LM² = xLM² + yLM² = 1² + 1² = 2 LM = √2 Distance AB xAB = xB - xA = 2 - 0 = 2 yAB = yB - yA = 2 - 0 = 2 AB² = xAB² + yAB² = 2² + 2² = 8 AB = √8 AB = 2√2 LM/AB = (√2) / (2√2) LM/AB = 1/2 → LM = (1/2).AB

OpenStudy (natasha18):

I'm not really smart enough to figure that ^^ out lol my teacher helped me :D

OpenStudy (whpalmer4):

well, you get a medal for all the typing, and getting your teacher to help you help someone else 😀

OpenStudy (natasha18):

Well then, thank you :)

OpenStudy (natasha18):

Did I help you? It was kind of confusing for me to understand lol

OpenStudy (whpalmer4):

yes, you helped me by saving me the effort of explaining the rest of the problem. I can't speak for the person who posted the question, however...

OpenStudy (natasha18):

@whpalmer4 I'm glad I could help :)